# HyGen: Efficient LLM Serving via Elastic Online-Offline Request Co-location

### v2

@@ 4.3 SLO-Aware Offline Scheduling Policies
After scheduling online requests, our offline scheduling policy repurposes the residual capacity to maximize throughput while ensuring fairness Sheng et al. (2024b).
To further optimize the serving throughput for offline requests, HyGen employs an SLO-aware Prefix Sharing Maximization (PSM) strategy. Prefix sharing is a widely-adopted technique for reusing the KV cache of shared input prefix between requests Gim et al. (2024); Zheng et al. (2024a); Zhao et al. (2024); Zheng et al. (2024b).
@@
Prefix Sharing Maximization Strategy.
Our Prefix Sharing Maximization (PSM) strategy organizes offline requests into a prefix tree that follows the structure of a Trie tree, capturing prefix sharing characteristics of all offline requests, with each leaf node representing an offline request. The priority of each request is determined by the Depth-First Search (DFS) order of the prefix tree, where requests with the greatest prefix sharing potential are scheduled together, maximizing prefix cache reuse among requests. Subsequently, HyGen performs SLO-aware offline scheduling (Section 4.1) using this order to maximize prefix sharing, reducing redundant computation and improving throughput. Additionally, the prefix tree structure ensures fast insertion of new arrivals and deletion of scheduled requests, guaranteeing efficient scheduling during runtime. A formalized algorithm is demonstrated in Appendix A.2. For example, consider a system that can process two offline requests per batch with the following request queue: (What is ML, How to code, What is AI, How to debug). Under traditional FCFS scheduling, requests are processed in arrival order: (What is ML, How to code), (What is AI, How to debug), resulting in no prefix sharing opportunities. In contrast, PSM’s prefix-aware scheduling reorders requests as: (What is ML, What is AI), (How to code, How to debug), enabling KV-cache reuse through shared prefixes.
The PSM strategy demonstrates strong extendability. Under certain scenarios, the vanilla PSM strategy may lead to starvation for requests with minimal prefix-sharing potential, blocked by continuous arrivals for an extended period. To illustrate this challenge, consider a request queue: (What is ML, What is AI, How to code, What is DL). When new requests arrive with similar prefixes (What is LLM, What is DNN), a naive prefix-sharing policy would continuously prioritize requests sharing the What is prefix, potentially starving the How to code request indefinitely. This issue can be mitigated by an extended version of our PSM policy, combining maximum prefix sharing with request freshness by using a utility ratio to ensure a balance between efficiency and fairness. Based on the utility ratio, a new offline request would be selected from the DFS order of the prefix tree or the most stale request from a self-balanced binary search tree sorted by freshness. A detailed algorithm is in Appendix A.3. These enhancements can make the PSM strategy more practical for real-world deployments, retaining its efficiency while improving fairness and adaptability.
5 Performance Evaluation
@@
5.1 Evaluation Setup
Implementation and Testbeds.
We implement HyGen on top of vLLM Kwon et al. (2023); vLLM Project (2023) and Sarathi-Serve Agrawal et al. (2024); Sarathi-Serve Project (2024), with 1,300 lines of additional code. We evaluate HyGen on two server configurations: one with 4 NVIDIA A100 GPUs (40GB VRAM each) and the other with 4 NVIDIA A40 GPUs (48GB VRAM each). Both servers have 64 CPU cores, 256GB DDR4 RAM, and a 1.5TB NVMe SSD.
Models and Workloads.
We evaluated HyGen using Llama2-7B Touvron et al. (2023) and Qwen-14B Bai et al. (2023) models on A100 and A40 GPUs, respectively.
For the online workload, we used the conversation trace from Azure LLM inference trace 2023 Patel et al. (2024), a one-hour production trace with real-world requests and timestamps. We randomly sampled the trace to achieve the desired QPS that suits our hardware serving capacity. Specifically, within a time duration of $T$ seconds, we would sample $T\times Q$ requests to suit a desired QPS $Q$. For the offline workload, we used arXiv summarization Cohan et al. (2018), a dataset for long document summarization.
Metrics.
For interference evaluation, we focus on both Time to First Token (TTFT) and Time Between Tokens (TBT) and their respective mean and 99th percentile (P99) values. We measure throughput using tokens per second (TPS) and queries per second (QPS).
Baselines.
For pure online inference, we use Sarathi Agrawal et al. (2024) as our baseline. For hybrid serving, we compare HyGen with two baselines for interference and throughput evaluation, respectively: (1) Sarathi++: We implement our online-first scheduling policy on Sarathi to support hybrid serving, including the request management and preemption handling policies introduced in Section 4.1. (2) HyGen*: To evaluate the throughput benefit of our HyGen design, we further improve Sarathi++ to an SLO-aware serving system, HyGen*. Besides inheriting the serving policies from Sarathi++, HyGen* serves offline requests at a specific offline QPS to control overall interference. The offline QPS is profiled using a similar design with the HyGen profiler to guarantee bounded SLO interference.
@@ 5.1 Evaluation Setup
Figure 6: Prefix Sharing Maximization policy achieves higher serving capacity.
Figure 7: SLO-aware profiler bridges the gap between batch latency and overall SLO.
Figure 8: Hygen dynamically controls offline throughput according to online workload and latency budget.
@@
Figure 12: Hygen meets SLOs and achieves higher throughput for Mooncake trace and Mistral-7B model.
@@
Figure 13: Hygen meets SLOs and achieves higher throughput.
@@ 5.2 End-to-end Performance
HyGen respects latency requirements in co-locating requests. We tested HyGen on four SLO metrics, mean TBT, P99 TBT, mean TTFT, and P99 TTFT. For each metric, we set varying interference tolerance ratios as the SLO. Figure 3 shows that HyGen is able to control the interference and guarantees to meet specific SLOs across our settings. Compared with Sarathi++, an SLO-unaware system that yields the same result for all metrics and tolerance ratios, HyGen showcases efficient SLO-aware latency control.
HyGen improves serving throughput. Figure 4 shows the offline throughput of HyGen for various metrics and tolerance ratios. Through efficient request co-location, HyGen improves overall serving throughput by up to 3.87$\times$ compared to pure online serving. Under the same SLO, HyGen consistently achieves higher throughput compared to HyGen*, yielding up to 5.84×5.84$\times$5.84 × offline throughput gain. This verifies that our fine-grained latency predictor and SLO-aware profiler designs achieve higher serving efficiency compared to their simplified counterparts.
@@ 5.3 Performance Breakdown
To test HyGen’s compatibility with prefix sharing, we conducted a simulation experiment using Azure LLM Inference as the online trace and MMLU Hendrycks et al. (2021a; b) as the offline dataset on a Llama2-7B model. In our simulation, we deducted the shared prompt prefix length between consecutive offline requests to simulate prefix sharing. Figure 7 shows that HyGen yields up to 4×4$\times$4 × offline throughput gain with its prefix sharing maximization scheduling policy.
@@ 5.4 Ablation Studies
To evaluate HyGen’s effectiveness in distributed inference, we deployed the Yi-34B model Young et al. (2024) on a server with 4 NVIDIA A40 GPUs using tensor-parallelism (TP) Shoeybi et al. (2019) and pipeline-parallelism (PP) Athlur et al. (2022); Narayanan et al. (2019); Huang et al. (2019) with degree 2 for each dimension. Using Azure LLM Inference and arXiv summarization workloads, Figure 9 shows that HyGen maintains its ability to meet SLOs and achieves higher offline throughput (up to 1.89×1.89$\times$1.89 ×) than the baseline.
@@
We further evaluate HyGen’s adaptability on two more experiments: The first using Mistral-7B model with Mooncake trace Qin et al. (2024), a trace containing request lengths and timestamps taken from real-world servers, as the online trace, and arXiv summarization as the offline trace; The second experiment uses Llama2-7B model with Azure LLM Inference trace and replaced the offline dataset with CNN/DailyMail summarization dataset See et al. (2017); Hermann et al. (2015). An analysis of the Mooncake trace is shown in Appendix B. Figure 12 and Figure 13 show that HyGen achieves superior performance than its counterparts under these settings.
@@
We tested HyGen’s robustness using several pre-trained LR latency predictors with varying prediction accuracy taken from other workloads and tested them on Azure LLM Inference trace and arXiv summarization dataset. Figure 15 shows how predictor accuracy (measured in mean absolute percentage error) affects offline throughput under the same P99 TBT SLO. HyGen remains robust across different accuracy settings. Additionally, our LR-based latency predictor is lightweight for training, with only $\sim$15ms training time for over 80,000 samples on CPUs.
@@
Under a high online arrival rate, online requests may approach system serving capacity, limiting the headroom for request co-location. Figure 15 shows the effect of online QPS on offline throughput with 5% P99 TBT tolerance. As online load increases, HyGen adjusts offline throughput based on the system’s residual serving capacity while maintaining higher throughputs.
@@ 6 Conclusion
This paper introduces HyGen, an LLM serving system that enables efficient co-location of online and offline workloads. We employ control mechanisms to predict and manage interference impacts, and a scheduling policy to opportunistically schedule offline serving. Evaluation on production workloads demonstrates that HyGen improves serving throughput by 3.87-5.84$\times$ while maintaining strict latency SLOs.
@@ A.1 HyGen Two-phase Scheduling Algorithm
16:        end if
17:        $m\leftarrow m-M_{off}$
18:        $B\leftarrow B\cup$ SLO_AWARE_SCHEDULE($R_{off}$, $Q_{off}$, $t$, $c$, $m$)
19:        $Q_{recv}$.send($B$, $t$, $m$)
20:     end while
21:  end function
22:  function SCHEDULER
23:     Output: batched requests $B$
24:     scheduled requests $B$, latency budget $t$, memory budget $m$leftarrow $Q_{recv}$.get(block=True)
25:     if REQUEST_ARRIVAL then
26:        UPDATE_BUDGET($B$, $t$, $m$)
27:        if $t<0$ or $m<0$ then
28:           PREEMPT_UNTIL_FIT($B$, $t$, $m$)
29:        end if
30:     end if
31:     $Q_{send}$.send($B$)
32:     return $B$
33:  end function
@@ A.2 SLO-Aware Prefix Sharing Maximization Algorithm
This section details the SLO-aware prefix sharing maximization algorithm design in HyGen. We construct a prefix tree $T_{p}$ to capture prefix sharing charateristics among offline requests. During scheduling, offline requests are selected in the DFS order of the prefix tree, and deleted once being scheduled. Additionally, running requests keep their original DFS order in future scheduling process, effectively utilizing prefix sharing.
@@
11:           end if
12:           $t\leftarrow t-t_{req}$
13:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
14:        else if $r$.state == PREFILL then
15:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
16:           if $l>0$ then
17:              $t\leftarrow t-t_{req}$
18:              $c\leftarrow c-l$
19:              $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
20:              $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
21:           else
22:              break
23:           end if
24:        end if
25:     end for
26:     while $T_{p}$ is not empty do
27:        $r\leftarrow T_{p}$.get_next_request()
28:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
29:        if $l>0$ then
30:           $t\leftarrow t-t_{req}$
31:           $c\leftarrow c-l$
32:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
33:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
34:           $T_{p}$.remove_request($r$)
35:        else
36:           break
37:        end if
38:     end while
39:     return $B$
40:  end function
@@ A.3 Extended SLO-Aware Prefix Sharing Maximization Algorithm
11:           end if
12:           $t\leftarrow t-t_{req}$, $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
13:        else if $r$.state == PREFILL then
14:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
15:           if $l>0$ then
16:              $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
17:           else
18:              break
19:           end if
20:        end if
21:     end for
22:     while $T_{p}$ is not empty do
23:        $rand\leftarrow$ RANDOM_NUMBER(0, 1)
24:        if $rand<u$ then
25:           $r\leftarrow T_{p}$.get_next_request()
26:        else
27:           $r\leftarrow T_{f}$.get_next_request()
28:        end if
29:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
30:        if $l>0$ then
31:           $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
32:           $T_{p}$.remove_request($r$)
33:           $T_{f}$.remove_request($r$)
34:        else
35:           break
36:        end if
37:     end while
38:     return $B$
39:  end function
Appendix B Mooncake Trace Analysis
This section provides an analysis of Mooncake trace, a production trace taken from real-world server with request lengths and request arrival timestamps. Figure 16 shows the varying request arrival rates of Mooncake trace over one-hour and ten-minute periods. This analysis further demonstrates the fluctuant characteristic of LLM services.
(a) 1 hour
(b) 10 minutes

### v4

@@ 4.3 SLO-aware Offline Scheduling Policies
After scheduling online requests, our offline scheduling policy repurposes the residual capacity to maximize throughput while ensuring fairness sheng2024fairness . To further optimize the serving throughput for offline requests, HyGen employs an SLO-aware Prefix Sharing Maximization (PSM) strategy. Prefix sharing is a widely adopted technique for reusing the KV cache of shared input prefixes between requests gim2024prompt ; zhao2024blendserve ; zheng2024sglang ; zheng2024batchllm .
Prefix Sharing Maximization Strategy.
Our PSM strategy organizes offline requests into a prefix tree following the structure of a Trie tree with each leaf node representing a request, capturing prefix sharing characteristics of all offline requests. The priority of each request is determined by the Depth-First Search (DFS) order of the prefix tree, where requests with the greatest prefix sharing potential are scheduled together. Subsequently, HyGen performs SLO-aware offline scheduling (Section 4.1) using this order to maximize prefix cache reuse, reducing redundant computation and improving throughput. The prefix tree structure also ensures fast insertion and deletion in runtime scheduling. A formalized algorithm is in Appendix A.2. For example, consider a system that can process two offline requests per batch with the following request queue: (What is ML, How to code, What is AI, How to debug). Under traditional FCFS scheduling, requests are processed in arrival order: (What is ML, How to code), (What is AI, How to debug), resulting in no prefix sharing opportunities. In contrast, PSM’s prefix-aware scheduling reorders requests as: (What is ML, What is AI), (How to code, How to debug), enabling KV cache reuse through shared prefixes.
The PSM strategy demonstrates strong extendability. Under certain scenarios, the vanilla PSM strategy may lead to starvation for requests with minimal prefix-sharing potential. Consider a request queue: (What is ML, What is AI, How to code, What is DL). When new requests arrive with similar prefixes (What is LLM, What is DNN), a naive prefix-sharing policy would continuously prioritize requests sharing the What is prefix, potentially starving the How to code request indefinitely. This issue can be mitigated by an extended version of our PSM policy, combining maximum prefix sharing with request freshness by using a utility ratio to ensure a balance between efficiency and fairness. Based on the utility ratio, a new offline request would be selected from the DFS order of the prefix tree or the most stale request from a self-balanced binary search tree sorted by freshness. A detailed algorithm is in Appendix A.3. These enhancements can make the PSM strategy more practical for real-world deployments, retaining its efficiency while improving fairness and adaptability.
5 Performance Evaluation
@@
Figure 6: Prefix Sharing Maximization improves serving throughput.
Figure 7: SLO-aware profiler contributes to HyGen’s improvements.
Figure 8: HyGen dynamically controls throughput according to online workload.
@@
Figure 12: HyGen meets SLOs and achieves higher throughput with CNN/DailyMail offline dataset.
(a) 1 hour
(b) 10 minutes
@@
Figure 14: HyGen meets SLOs and achieves higher throughput for Mooncake trace.
(a) SLOs
(b) Throughput
Figure 15: HyGen meets SLOs and achieves higher throughput on A5000 GPU and Sheared-LLaMA-2.7B model.
@@
5.1 Evaluation Setup
Implementation and Testbeds.
We implement HyGen on top of vLLM kwon2023efficient ; vllm2023github and Sarathi sarathi ; sarathi2024github , with 1,300 lines of additional code. We evaluate HyGen on three server configurations: one with 4 NVIDIA A100 GPUs (40GB VRAM each), one with 4 NVIDIA A40 GPUs (48GB VRAM each), and one with 1 NVIDIA A5000 GPU (24GB VRAM). All servers have 64 CPU cores, 256GB DDR4 RAM, and a 1.5TB NVMe SSD.
Models and Workloads.
For end-to-end evaluation, we use Llama2-7B touvron2023llama and Qwen-14B bai2023qwen models on A100 and A40 GPUs, respectively. Online workloads are based on the conversation trace from Azure LLM inference trace 2023 patel2024splitwise , a one-hour production trace with real-world requests and timestamps. We randomly sampled the trace to achieve the desired QPS that suits our hardware serving capacity. Specifically, within a time duration of $T$ seconds, we would sample $T\times Q$ requests to suit a desired QPS $Q$. For offline workloads, we use arXiv summarization cohan-etal-2018-discourse , a dataset for long document summarization. In our ablation studies, we evaluate HyGen across different model scales ranging from Sheared-LLaMA-2.7B xia2023sheared , Mistral-7B jiang2023mistral7b to Yi-34B young2024yi . The Mooncake trace mooncake is further used as the online trace, providing industrial request length distributions and arrival patterns, while CNN/DailyMail hermann2015teaching ; see-etal-2017-get and MMLU hendrycks2021ethics ; hendryckstest2021 are used as offline traces in the ablation studies. For interference evaluation, we focus on Time to First Token (TTFT) and Time Between Tokens (TBT), including their mean and 99th percentile (P99) values. Throughput is measured in tokens per second (TPS) and queries per second (QPS).
Baselines.
For pure online inference, we use Sarathi sarathi as our baseline. For pure offline serving, we use Sarathi-offline to evaluate the maximum offline serving capacity, where an optimal chunk size is profiled for offline workload to maximize throughput. The hyperparameter search of Sarathi-offline achieves $\sim$12% throughput gain compared to the default setup, ensuring optimal baseline performance for fair comparison. We then compare HyGen with two baselines for interference and throughput evaluation, respectively: (1) Sarathi++: We implement our online-first scheduling policy on Sarathi to support hybrid serving, including the request management and preemption handling policies introduced in Section 4.1. (2) HyGen*: To evaluate the throughput benefit of our HyGen design, we further improve Sarathi++ to an SLO-aware serving system, HyGen*. Besides inheriting the serving policies from Sarathi++, HyGen* serves offline requests at a specific offline QPS to control overall interference. The offline QPS is profiled using a similar design with the HyGen profiler to guarantee bounded SLO interference.
@@ 5.2 End-to-end Performance
HyGen respects latency requirements in co-locating requests. We evaluate HyGen under four SLO metrics (mean TBT, P99 TBT, mean TTFT, and P99 TTFT) with varied interference tolerance ratios. Figure 3 shows that HyGen controls interference and guarantees to meet specific SLOs across our settings. Compared with Sarathi++, an SLO-unaware system that yields the same result for all metrics and tolerance ratios, HyGen shows efficient SLO-aware latency control.
[…]
@@ 5.3 Performance Breakdown
To test HyGen’s compatibility with prefix sharing, we conducted a simulation experiment using Azure LLM Inference as the online trace and MMLU hendrycks2021ethics ; hendryckstest2021 as the offline dataset on a Llama2-7B model. In our simulation, we deducted the shared prompt prefix length between consecutive offline requests to simulate prefix sharing. Figure 7 shows that HyGen yields up to $4\times$ offline throughput gain with its prefix sharing maximization scheduling policy.
@@ 5.4 Ablation Studies
To evaluate HyGen’s effectiveness in distributed inference, we deployed the Yi-34B model young2024yi on a server with 4 NVIDIA A40 GPUs using tensor-parallelism (TP) shoeybi2019megatron and pipeline-parallelism (PP) athlur2022varuna ; huang2019gpipe ; narayanan2019pipedream with degree 2 for each dimension. Using Azure LLM Inference and arXiv summarization workloads, Figure 9 shows that HyGen maintains its ability to meet SLOs and achieves higher offline throughput (up to $1.89\times$) than the baseline.
@@
We further evaluate HyGen’s adaptability on two more experiments: The first using Mistral-7B model jiang2023mistral7b with Mooncake trace mooncake , a trace containing request lengths and timestamps taken from real-world servers, as the online trace, and arXiv summarization as the offline trace; The second experiment uses Llama2-7B model with Azure LLM Inference trace and replaced the offline dataset with CNN/DailyMail summarization dataset hermann2015teaching ; see-etal-2017-get . Figure 14 shows the varying request arrival rates of Mooncake trace over one-hour and ten-minute periods, further demonstrating the fluctuating nature of LLM services. Figure 12 and Figure 14 show that HyGen achieves superior performance than its counterparts under these settings.
Impact of hardware testbeds.
[…]
@@
[…]
@@
Figure 17 shows the effect of online QPS on offline throughput with 5% P99 TBT tolerance. As online load increases, HyGen adjusts offline throughput based on the system’s residual serving capacity while maintaining higher throughputs. Understandably, a high online arrival rate limits the headroom for co-location as it approaches system serving capacity.
@@ 6 Conclusion
[…]
Acknowledgements
We thank the anonymous reviewers for their constructive and insightful feedback. This work was supported in part by grants from Cisco and Google, and by an award from NVIDIA Academic Program. It also utilized the Delta system at the National Center for Supercomputing Applications (NCSA) through allocation CIS240236 from the ACCESS program.
@@ A.1 HyGen Two-phase Scheduling Algorithm
16:   $m\leftarrow m-M_{off}$
17:   $B\leftarrow B\cup$ SLO_AWARE_SCHEDULE($R_{off}$, $Q_{off}$, $t$, $c$, $m$)
18:   $Q_{recv}$.send($B$, $t$, $m$)
19:function SCHEDULER
20:  Output: batched requests $B$
21:  scheduled requests $B$, latency budget $t$, memory budget $m$ $\leftarrow$ $Q_{recv}$.get(block=True)
22:  if REQUEST_ARRIVAL then
23:   UPDATE_BUDGET($B$, $t$, $m$)
24:   if $t<0$ or $m<0$ then
25:    PREEMPT_UNTIL_FIT($B$, $t$, $m$)
26:  $Q_{send}$.send($B$)
27:  return $B$
@@ A.2 SLO-Aware Prefix Sharing Maximization Algorithm
This section details the SLO-aware prefix sharing maximization algorithm design in HyGen. We construct a prefix tree $T_{p}$ to capture prefix sharing characteristics among offline requests. During scheduling, offline requests are selected in the DFS order of the prefix tree, and deleted once being scheduled. Additionally, running requests keep their original DFS order in future scheduling process, effectively utilizing prefix sharing.
@@
11:    $t\leftarrow t-t_{req}$
12:    $B\leftarrow B\cup$ {($r$, 0, $t_{req}$)}
13:   else if $r$.state == PREFILL then
14:    $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
15:    if $l>0$ then
16:     $t\leftarrow t-t_{req}$
17:     $c\leftarrow c-l$
18:     $m\leftarrow m$- GET_NUM_BLOCKS($l$)
19:     $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
20:    else
21:     break
22:  while $T_{p}$ is not empty do
23:   $r\leftarrow T_{p}$.get_next_request()
24:   $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
25:   if $l>0$ then
26:    $t\leftarrow t-t_{req}$
27:    $c\leftarrow c-l$
28:    $m\leftarrow m$- GET_NUM_BLOCKS($l$)
29:    $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
30:    $T_{p}$.remove_request($r$)
31:   else
32:    break
33:  return $B$
@@ A.3 Extended SLO-Aware Prefix Sharing Maximization Algorithm
11:    $t\leftarrow t-t_{req}$, $B\leftarrow B\cup$ {($r$, 0, $t_{req}$)}
12:   else if $r$.state == PREFILL then
13:    $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
14:    if $l>0$ then
15:     $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m$- GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
16:    else
17:     break
18:  while $T_{p}$ is not empty do
19:   $rand\leftarrow$ RANDOM_NUMBER(0, 1)
20:   if $rand<u$ then
21:    $r\leftarrow T_{p}$.get_next_request()
22:   else
23:    $r\leftarrow T_{f}$.get_next_request()
24:   $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
25:   if $l>0$ then
26:    $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m$- GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
27:    $T_{p}$.remove_request($r$)
28:    $T_{f}$.remove_request($r$)
29:   else
30:    break
31:  return $B$
A.4 Complexity Analysis of the Two-Phase Scheduling Algorithm
In this section, we present an analysis of the computational complexity of HyGen’s two-phase scheduler. The time complexity of its core components is as follows:
  • Latency Prediction: $O(1)$ inference using a pre-trained LR model.
  • PSM: The initial construction of the prefix tree is $O(N×L)$, where $N$ is the number of offline requests and $L$ is the average number of tokens. Each insertion or deletion costs $O(L)$. In implementation, getting the next request in the DFS order only takes $O(1)$, since the DFS order is put in a pre-processed list derived from the prefix tree and can be synced up with the prefix tree asynchronously.
  • PSM with Fairness: In the fairness-aware PSM algorithm, a self-balancing BST is used for picking the stalest request. Each lookup, insertion, or deletion takes $O(\log n)$. In implementation, the requests can be kept in an FCFS queue, which syncs up with the BST asynchronously to guarantee correctness, so that each lookup still only takes $O(1)$ time.
Overall, for the two-phase scheduling using existing policies, the time complexity remains $O(n)$ as asynchronous updates can be performed for advanced policies like fairness-aware PSM.
Appendix B Further Discussion of the Latency Predictor
Effective co-location in HyGen hinges on a latency predictor that is both highly accurate and computationally lightweight, as it enables real-time scheduling decisions without introducing significant overhead. To this end, we employ a linear regression (LR) model, which provides inference in constant time, $O(1)$, making it ideal for a dynamic serving environment. This section details the model’s formulation and its adaptability to the complexities of real-world deployment scenarios.
Model Formulation.
The execution time of a serving batch, $T_{\text{batch}}$, is primarily determined by the computational patterns of its prefill and decode stages. We model this relationship as a function of the batch’s composition:
$T_{\text{batch}}=f(S_{p},S_{d},S_{p}^{2},N_{p},N_{d})$ (2)
where $S_{p}$ and $S_{d}$ represent the total number of tokens in the prefill and decode phases, respectively, and $N_{p}$ and $N_{d}$ are the corresponding request counts. The quadratic term $S_{p}^{2}$ is crucial for capturing the non-linear scaling of the self-attention mechanism, which dominates the computational cost of the prefill stage. In contrast, the decode stage, which processes one token per request at a time, exhibits linear scaling with the batch size ($N_{d}$). The model’s coefficients are pre-trained on data gathered by systematically profiling the target hardware and LLM across a diverse set of batch compositions, a lightweight process that ensures applicability to any deployment environment.
Robustness Through System-Level Design.
While a linear model offers unparalleled efficiency, its accuracy in the face of dynamic system conditions (e.g., GPU temperature variations, resource contention) is a critical consideration. HyGen ensures robustness not through a more complex model, but through a synergistic system design. The SLO-aware profiler (Section 4.2) first establishes a macro-level operational latency budget, which implicitly captures the system’s current performance characteristics. The LR predictor then operates at a micro-level, making fine-grained decisions on batch composition within this pre-calibrated budget.
This two-level approach effectively decouples the real-time scheduling decision from low-level hardware variability. As shown in Figure 17, this design allows the system to maintain robust performance and meet SLOs even in scenarios with predictor error rates exceeding 20%. Furthermore, the feature set is extensible; environmental factors like hardware load can be readily integrated into the model if required for specific use cases.
Adaptability to Modern Model Architectures.
The feature set defined in Equation 2 is sufficiently general to generalize across various modern LLM architectures without modification. For instance, for Mixture-of-Experts (MoE) models [48], where a fixed number of experts are activated per token, the resulting computational cost scales linearly with the total number of tokens processed and is effectively modeled by the $S_{p}$ and $S_{d}$ features. Similarly, in hybrid architectures that combine linear-complexity components (e.g., Mamba [19]) with quadratic-complexity attention, our model naturally captures both computational patterns; the linear cost is reflected in the learned coefficient for $S_{p}$, while the quadratic cost of the Transformer blocks is captured by the $S_{p}^{2}$ term. This adaptability underscores the predictor’s robust design, ensuring its relevance as LLM architectures continue to evolve.
Appendix C Taming the Throughput-Latency Tradeoff in LLM Serving with HyGen
Traditional instance scaling solutions address bursty workloads by launching new instances, which can cause tens of seconds to several minutes cold-start delays [15, 64, 66]. To handle these delays, providers often keep standby instances online, leading to wasted resources during off-peak periods.
In contrast, HyGen optimizes resource utilization by running offline workloads on idle resources and reallocating them to online requests in real-time, within a single inference iteration. This eliminates cold-start delays while ensuring high resource utilization.
HyGen complements instance scaling solutions by automating the transition between online and offline workloads, reducing the need for manual intervention. While instance scaling manages large load fluctuations, HyGen ensures efficient resource use during low-traffic periods, optimizing overall system performance in fixed-size clusters.